# Single-temporal Supervised Remote Change Detection for Domain Generalization

### v3

@@ 2.1 Supervised Change Detection
RSCD detects the object changes by comparing the semantic features of pixels representing the same geographical region. Deep learning neural networks have demonstrated their excellence in change detection and these methods are based on siamese structure [28, 32, 20]. Existing supervised change detection methods often aim to improve model performance through network structure optimization or the incorporation of effective tricks, such as feature exchange [28, 32], attention mechanism [17, 18, 6, 21], feature enhancement [17]. These supervised methods heavily rely on the availability of extensive labeled datasets and require training a separate model for each dataset. It results in overfitting issues, as the models become too specialized to the dataset and the number of parameters rises. Consequently, the generalization performance of these models across different domains is significantly compromised.
@@ 2.3 Single-Temporal Supervised Change Detection
Existing high-quality annotated change detection datasets [5, 13, 15] are very small, but there are already a large number of remote sensing single-temporal annotated datasets [11, 31]. ChangeStar [38] first proposed a single-temporal change detection training strategy STAR [38]. However, it is too simple and crude that STAR forms the pseudo-pair by randomly matching within the training batch. It goes against the changes in the real world, significantly different from the true distribution. Moreover, it is difficult to perform shape analysis on the corresponding objects in the two randomly matched images.
In this paper, we propose a single-temporal training strategy based on Artificial Intelligence Generative Content (AIGC). By generating abundant trainable data through directed category editing, we mitigate data dependency. This approach constructs data closely resembling the natural real-world distribution, significantly enhancing model generalization.
@@ 3.4 Single-Temporal Controllable AI-Generated Training Strategy
Revisiting the essence of temporal image pairs for change detection, it is evident that pixels at the same location in these pairs represent two distinct states of the same geographical area at different times. These states may manifest as different physical conditions of the same object over time or as changes induced between different categories of objects at the same location. The establishment of correspondence between these diverse states is fundamental to train object change detection models based on single-temporal images. Acquiring high-quality instance annotations for such bi-temporal image pairs comes at a considerable cost. Nevertheless, a wealth of single-temporal remote sensing instance segmentation images is already available and highly relevant to change detection. ChangeStar [38] proposes a method of constructing image pairs by STAR [38]. However it is crude, deviates from practical significance, and violates the true laws observed in the natural world.

### v4

@@ 2.1 Supervised Change Detection
RSCD detects the object changes by comparing the semantic features of pixels representing the same geographical region. Deep learning neural networks have demonstrated their excellence in change detection and these methods are based on siamese structure [29, 33, 21]. Existing supervised change detection methods often aim to improve model performance through network structure optimization or the incorporation of effective tricks, such as feature exchange [29, 33], attention mechanism [17, 18, 6, 22], feature enhancement [17]. These supervised methods heavily rely on the availability of extensive labeled datasets and require training a separate model for each dataset. It results in overfitting issues, as the models become overly specialized to the dataset and the number of parameters rises. Consequently, the generalization performance of these models across different domains is significantly compromised.
@@ 2.3 Single-Temporal Supervised Change Detection
Existing high-quality annotated change detection datasets [5, 13, 15] are very small, but there are already a large number of remote sensing single-temporal annotated datasets [11, 32]. ChangeStar [39] first proposed a single-temporal change detection training strategy STAR [39] which forms the pseudo-pair by randomly matching within the training batch. However, it lacks changes over long temporal span and leads to a significant number of changes that do not correspond to reality. It goes against the changes in the real world, significantly different from the true distribution. Moreover, it is difficult to perform shape analysis on the corresponding objects in the two randomly matched images. Changen [40] generates pseudo-image pairs based GAN [19], but its controllability and flexibility are limited.
In this paper, we propose a single-temporal training strategy based on Artificial Intelligence Generative Content (AIGC) with text-image mask controllable generation. By generating abundant trainable data through directed category editing, we mitigate data dependency. This approach constructs data closely resembling the natural real-world distribution, significantly enhancing model generalization.
@@ 3.4 Single-Temporal Controllable AI-Generated Training Strategy
Revisiting the essence of temporal image pairs for change detection, it is evident that pixels at the same location in these pairs represent two distinct states of the same geographical area at different times. These states may manifest as different physical conditions of the same object over time or as changes induced between different categories of objects at the same location. The establishment of correspondence between these diverse states is fundamental to train object change detection models based on single-temporal images. Acquiring high-quality instance annotations for such bi-temporal image pairs comes at a considerable cost. Nevertheless, a wealth of single-temporal remote sensing instance segmentation images is already available and highly relevant to change detection. ChangeStar [39] proposes a method of constructing image pairs by STAR [39]. However there are some shortcomings, the generated data deviate from practical significance, and violate the true changes observed in the natural world.